Cody’s car accelerates from 0m/s to 45 m/s northward in 15 seconds. What is the acceleration of the car?



1
Expert's answer
2021-10-04T15:30:37-0400

By definition, the acceleration is given as follows:


"a = \dfrac{v_f-v_i}{t}"

where "v_i = 0m\/s, v_f = 45m\/s" is the initial and final speeds respectively, "t = 15s". Thus, obtain:


"a = \dfrac{45m/s-0m/s}{15s} = 3m/s^2"

Answer. 3 m/s^2.